OpenStudy (anonymous):

Let A and B be square matrices of the same size. If A and B are both ______, so is AB. 1) symmetric 2) invertible 4) diagonalisable 8) upper triangular

OpenStudy (anonymous):

I know 2 and 8 are correct while 1 is not. Not sure about 4.

OpenStudy (anonymous):

For 1) \[(AB)^T = B^T A^T = BA \ne AB \]

OpenStudy (experimentx):

that's correct

OpenStudy (anonymous):

4 is correct?

OpenStudy (experimentx):

hmm ... I have to think about it. what do you think about 2 and 8 ??

OpenStudy (experimentx):

most likely no for 4

OpenStudy (anonymous):

2 and 8 are correct, I think. for 2) (AB)^(-1) (AB)= B^-1 A^-1 (AB) = B^(-1) B = I = AB( B^-1 A^-1 ) = AB ( AB)^(-1)

OpenStudy (anonymous):

For 8, we can use block multiplication to prove it

OpenStudy (experimentx):

2 is 100% correct

OpenStudy (experimentx):

8 is also correct.

OpenStudy (experimentx):

about 4 ... do you mean Eiven value decomposition ?

OpenStudy (anonymous):

*block multiplication and induction for 8

OpenStudy (anonymous):

Hmm, not sure.

OpenStudy (experimentx):

could you check your book? most probably it's eiven value decomposition http://en.wikipedia.org/wiki/Diagonalizable_matrix#How_to_diagonalize_a_matrix

OpenStudy (anonymous):

Couldn't find it.... I actually tried this: \[A = PD_1P^{-1}, B=QD_2Q^{-1}\]Then,\[AB = PD_1P^{-1}QD_2Q^{-1}\] But this is ugly

OpenStudy (experimentx):

it was due to this factor I thought most likely no ... becuase if those two P^-1 and Q are permutation matrices ,, they would wind up the D's and the matrix would not be diagonaliable. But earlier ... loser66 mentioned that invertible matrix are diagonalizable. and it seems that it's true.

OpenStudy (anonymous):

A matrix is invertible does not imply that it is diagonalizable

OpenStudy (experimentx):

what the hell ... I thought you could eliminate using Gauss Jordan elimination.

OpenStudy (experimentx):

let's assume this is true ... and try to find a counter example.

OpenStudy (experimentx):

if P^-1Q = I, then it's 100% diagonalizable.

OpenStudy (experimentx):

that implies P=Q

OpenStudy (loser66):

Question: The statement is " if both A, B are diagonalizable, then so AB" It doesn't relate to invertible or not. Am I right?

OpenStudy (loser66):

so, it's true.

OpenStudy (anonymous):

Whether A or B is invertible does not guarantee AB is diagonalizable. Even if a matrix is invertible, it can be not diagonlizable. An counter example can be found in link I posted earlier

OpenStudy (experimentx):

I thought pivots are going to act as diagonal elements.

OpenStudy (loser66):



OpenStudy (anonymous):